Jack's probability of passing a math course is 40%, and Larry's probability of passing the same course is 70%. If the two events are independent, find the following probabilities.
a) P(both of them will pass statistics)
b) P(at least one of them will pass statistics)

Answer by FrankM(1040)

Both will pass is .4 * .7 = .28 or 28%
At least 1 will pass is 1-(both will fail) = 1-(.6*.3)= .82 or 82%